# Smooth Lower Bounds for Differentially Private Algorithms via Padding-and-Permuting Fingerprinting Codes

### v3

@@ 2.3 More General Framework
2.4 Comparison with [NME22]
[NME22, ] developed very different hard-instances for lower bounding user-level DP averaging that can be used to prove a similar statement to Theorem 1.3 (our DP averaging lower bound). In their setting, a distribution vector $p=(p_{1},\ldots,p_{d})\in[0.5/d,1.5/d]^{d}$, $\sum_{i=1}^{d}p_{i}=1$, is sampled, and the goal is to estimate it. For each user (out of $m$), they sample $m=\lambda^{2}d$ one-hot vectors according to $p$, and provide the average of the vectors as the user’s input point. Since each $p_{i}\approx 1/d$, it can be shown that w.h.p., the resulting points have $\ell_{2}$ diameter $\gamma\approx 1/\sqrt{m}=1/\lambda\sqrt{d}$ (and also close to $p$ up to such an additive error). So their Lemma 23 essentially states that estimating $p$ up-to additive $\ell_{2}$ error of $1/\sqrt{d}=\lambda\cdot\gamma$, requires $\tilde{\Omega}(\sqrt{d}/\lambda)$ users (which matches our Theorem 1.3 up to the low-order terms).
While both works yield similar smooth lower bound for DP averaging, the focus of the works is different: [NME22] focus on user-level DP averaging, while our focus is on general bounds for various (item-level) DP problems. As part of that, we have several advantages in the item-level DP case:
@@ 2.4 Comparison with [NME22]
    Tighter Bounds: [NME22] have an $1/\log^{7}(d/\lambda)$ dependency in their DP-averaging lower bound (see their Theorem 10), while we only have an $1/\log^{1.5}(d/\lambda)$ dependency.
@@
    Boolean instances: We create boolean hard-instances, while [NME22]’s instances are not.
@@ 4 Strong Fingerprinting Lemma
In this section we adopt the notation of [DSS+15, ]. For $p\in[-1,1]$ let $x\sim p$ denote that $x\in\{\pm 1\}$ is drawn with ${\mathrm{E}}\mathopen{}\mathclose{{}\left[x}\right]=p$, and let $x_{1\ldots n}\sim p$ denote that each $x_{i}$ is sampled independently with ${\mathrm{E}}\mathopen{}\mathclose{{}\left[x_{i}}\right]=p$ (that is, ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[x_{i}=1}\right]=\frac{1+p}{2}$ and ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[x_{i}=-1}\right]=\frac{1-p}{2}$).
@@ 4.1 One Dimension
where we used the inequality $e^{-1.1y}\leq 1-y$ for every $y\in[0,1/6]$. Hence, we conclude that
@@ 4.2 Increasing the Dimension
The first part of the lemma follows from similar arguments as in the proof of Proposition 10 in [DSS+15]. In particular, for a given $j\in[d]$, we fix all columns except for the $j^{\rm th}$ column. Since the mechanism $\mathsf{M}$ only gets $\mathbb{X}$ as an input, and each columns $j$ of $\mathbb{X}$ is sampled independently, then there is some $F_{j}\colon\{-1,1\}^{n}\rightarrow[-1,1]^{d}$ for which $\mathbb{q}^{j}\sim F_{j}(\mathbb{x}^{j})$, where $\mathbb{x}^{j}=(\mathbb{x}_{1}^{j},\dots,\mathbb{x}_{n}^{j})$ (the $j$’th column of $\mathbb{X}$), such that Section 4.1 holds for it. Then, by this and the fact that $\mathsf{M}$ is a strongly-accurate mechanism,
@@
The second part is proven similarly to Lemma 11 in [DSS+15]. Let us define the random variable $\mathbb{a}_{j}=\mathbb{q}^{j}\cdot\sum_{i\in[n]}(\mathbb{x}_{i}^{j}-\mathbb{z}^{j})$ for every $j\in[d]$. Then, we get that
@@
and from the first part we have ${\mathrm{E}}\mathopen{}\mathclose{{}\left[\sum_{j=1}^{d}\mathbb{a}_{j}}\right]\geq\frac{0.4d}{\ln(5n)}$. In order to use Hoeffding’s inequality on the sum $\sum_{j=1}^{d}\mathbb{a}_{j}$, we first assume that the random variables $\mathbb{a}_{1},\dots,\mathbb{a}_{d}$ are independent. Later, as in [DSS+15], we show how to remove this assumption. Next, observe that $-2n\leq\mathopen{}\mathclose{{}\left|\mathbb{a}_{j}}\right|\leq 2n$ for every $j\in[n]$. Therefore, using an Hoeffding’s inequality (1), it follows that
@@ 5 Framework for Lower Bounds
Let $(\mathbb{x}_{1},\ldots,\mathbb{x}_{n_{0}},\mathbb{z})\sim{\cal D}(n_{0},d_{0})$, let $\mathbb{X}=(\mathbb{x}_{1},\ldots,\mathbb{x}_{n_{0}})$, and let $\mathbb{q}=\mathsf{A}(\mathbb{X})$. By Section 5 and the union bound, we deduce that there exists $i\in[n_{0}]$ such that
@@
In the following, let $\mathbb{p}_{1},\ldots,\mathbb{p}_{d_{0}}\in[-1,1]$ be the (random variables of the) expectations that were chosen as part of the sampling of $(\mathbb{x}_{1},\ldots,\mathbb{x}_{n_{0}},\mathbb{z})$ from ${\cal D}(n_{0},d_{0})$ (as defined in Section 4.2), and in the rest of the proof we fix $(\mathbb{p}_{1},\ldots,\mathbb{p}_{d_{0}})=(p_{1},\ldots,p_{d_{0}})$ such that Equation 4 holds under this fixing. For the above $i\in[n_{0}]$, let $\mathbb{x}_{i}^{\prime}\in\{-1,1\}^{d_{0}}$ be an independent random variable such that each $(\mathbb{x}_{i}^{\prime})^{j}\sim p_{j}$ (independently) for every $j\in[d_{0}]$. Note that $\mathbb{x}_{1},\ldots,\mathbb{x}_{n_{0}},\mathbb{x}_{i}^{\prime},\mathbb{z}$ are i.i.d. random variables. However, while $\mathbb{q}$ depends on $\mathbb{x}_{i}$, it is independent of $\mathbb{x}_{i}^{\prime}$ and $\mathbb{z}$. Assume towards a contradiction that $\mathsf{A}$ is $\mathopen{}\mathclose{{}\left(1,\frac{\beta}{4n_{0}}}\right)$-DP. By applying Section 3.4.1 with the random predicate $P(x_{i},q)=\mathds{1}_{\{\langle x_{i},q\rangle-\langle\mathbb{z},q\rangle>\frac{0.2d_{0}}{n_{0}\ln(5n_{0})}\}}$, we deduce from Equation 4 that
@@
In the following, we prove that since $\mathbb{x}_{i}^{\prime}$ is independent of $\mathbb{q}$, then the above probability is much smaller, which will lead to a contradiction. For every $j\in[d_{0}]$, define the random variable $\mathbb{b}_{j}=\mathbb{q}^{j}((\mathbb{x}_{i}^{\prime})^{j}-\mathbb{z}^{j})$. Then, we get that
@@ 6 Padding And Permuting (PAP) FPC
The first inequality hold by Section 6. The last inequality holds since $\mathsf{M}$ is $(\alpha,\beta)$-weakly-accurate and for every $X\in\{-1,1\}^{n\times d_{0}}$ and $Y\in\operatorname{Supp}(Y_{X})$ it holds that $\mathopen{}\mathclose{{}\left|{\cal J}_{Y}^{1}}\right|,\mathopen{}\mathclose{{}\left|{\cal J}_{Y}^{-1}}\right|\geq\ell\geq\frac{1}{2}(1-\alpha)d$. Thus by Section 5, $n\geq\Omega(\sqrt{d_{0}}/\log^{1.5}(d_{0}/\beta))$, and the proof follows since $d_{0}=\Theta(\alpha d)$.     $\Box$
@@
We proof Theorem 1.2 using multiple PAP-FPC copies.
@@
    For $t\in[k]$ set $Y_{t}=\begin{cases}Y&t=s\\ \mathsf{PAP}(A_{t},P_{t})&t\neq s\end{cases}$. Denote $Y_{t}=(x_{(t-1)n_{0}+1}^{\prime},\>\ldots,\>x_{tn_{0}}^{\prime})\in\{-1,1\}^{n_{0}\times d}$.
@@
Define $\mathsf{G}\colon\{-1,1\}^{n_{0}\times d_{0}}\times{\cal V}\rightarrow\{-1,1\}^{n\times d}$ as the algorithm that on inputs $X\in\{-1,1\}^{n_{0}\times d_{0}}$ and $v=(s,\vec{P})\in{\cal V}$ for $s\in[k]$ and $\vec{P}=(P_{1},\ldots,P_{k})\in({\cal P}_{d})^{k}$: Computes $Y=\mathsf{PAP}(X,P_{s})$, and outputs $X^{\prime}\sim\mathsf{H}(Y,v)$. Note that for every $v\in{\cal V}$, $\mathsf{G}(\cdot,v)$ is neighboring-preserving (Section 3.4.1). Define $\mathsf{F}\colon{\cal V}\times[-1,1]^{d}\rightarrow[-1,1]^{d_{0}}$ as the algorithm that on inputs $v=(s,\vec{P})\in{\cal V}$ and $w\in[-1,1]^{d}$, outputs $(w\cdot P_{s}^{T})^{1,\ldots,d_{0}}$. Note that $\mathsf{F}$ is deterministic, and the random choice of $\mathsf{G}$ is $\vec{A}\sim{\cal D}^{k}$ (chosen when executing $\mathsf{H}$). Our goal is to show that
@@
We now focus on proving Equation 8. Consider the mechanism $\mathsf{M}^{\prime}\colon\{-1,1\}^{n_{0}\times d}\rightarrow[-1,1]^{d}$ that on input $Y\in\{-1,1\}^{n_{0}\times d}$, samples $v=(s,\vec{P})\leftarrow{\cal V}$, computes $X^{\prime}=\mathsf{H}(Y,v)$ and outputs $w\sim\mathsf{M}(X^{\prime})$. In the following, let $\mathbb{P}\leftarrow{\cal P}_{k}$ be a random variable, and for $X\in\{-1,1\}^{n_{0}\times d_{0}}$ define $\mathbb{Y}_{X}=\mathsf{PAP}(X,\mathbb{P})$. Compute
@@
The first equality holds since by the definitions of $\mathsf{M}^{\prime}$ and $\mathsf{G}$, for every $X$, the pair $\mathopen{}\mathclose{{}\left(\mathsf{M}^{\prime}(\mathbb{Y}_{X}),\mathbb{Y}_{X}}\right)$ has the same distribution as the pair $\mathopen{}\mathclose{{}\left(\mathsf{M}(\mathsf{G}_{s,\vec{A}}(X,v)),\mathsf{PAP}(X,P_{s})}\right)_{v=(s,\vec{P})\leftarrow{\cal V},\>\vec{A}\sim{\cal D}^{k}}$. The second equality holds since $\mathsf{G}_{\vec{A}}(\cdot,(s,\vec{P}))$ does not use the value of $A_{s}$ which is drawn as $X$. The third equality holds by the definition of $\mathsf{G}$. The inequality holds since $\mathsf{M}$ is an $(k,\alpha,\beta)$-weakly-accurate mechanism, and it gets $X^{\prime}=(x_{1}^{\prime},\ldots,x_{n}^{\prime})$ where each $Y_{t}=(x^{\prime}_{(t-1)n_{0}+1},\>\ldots,\>x^{\prime}_{tn_{0}})$ has $\mathopen{}\mathclose{{}\left|{\cal J}_{Y_{t}}^{1}}\right|,\mathopen{}\mathclose{{}\left|{\cal J}_{Y_{t}}^{-1}}\right|\geq\ell\geq\frac{1}{2}(1-\alpha)d$. Thus w.p. $\beta$, $\mathsf{M}$ is guaranteed to output $w\in[-1,1]^{d}$ that strongly-agrees with one of the $Y_{1},\ldots,Y_{k}$. But since they are drawn independently from the same distribution, the probability to agree with $Y_{s}$ for a random $s\leftarrow[k]$ decreases by a factor $k$, and overall is at least $\beta/k$. We now apply Section 6 to get that
@@ 7.1 Averaging
7.2 $k$-Means Clustering
In this section, we prove Theorem 1.4. We start by recalling the notations.
Let ${\cal B}_{d}:=\{x\in{\mathbb{R}}^{d}\colon\mathopen{}\mathclose{{}\left\lVert x}\right\rVert_{2}\leq 1\}$. For a database $S\in({\cal B}_{d})^{n}$ and $k$ centers $C=(c_{1},\ldots,c_{k})\in({\cal B}_{d})^{k}$, let
${\rm COST}_{S}(C):=\sum_{x\in S}\min_{i\in[k]}\mathopen{}\mathclose{{}\left\lVert x-c_{i}}\right\rVert_{2}^{2}\quad\text{ and }\quad{\rm OPT}_{k}(S):=\min_{C\in({\mathbb{R}}^{d})^{k}}{\rm COST}_{S}(C).$

### v4

@@ 2.3 More General Framework
2.4 Comparison with [NarayananME22]
[…]
NarayananME22 developed very different hard-instances for lower bounding user-level DP averaging that can be used to prove a similar statement to Theorem 1.3 (our DP averaging lower bound). In their setting, a distribution vector $p=(p_{1},\ldots,p_{d})\in[0.5/d,1.5/d]^{d}$, $\sum_{i=1}^{d}p_{i}=1$, is sampled, and the goal is to estimate it. For each user (out of $m$), they sample $m=\lambda^{2}d$ one-hot vectors according to $p$, and provide the average of the vectors as the user’s input point. Since each $p_{i}\approx 1/d$, it can be shown that w.h.p., the resulting points have $\ell_{2}$ diameter $\gamma\approx 1/\sqrt{m}=1/\lambda\sqrt{d}$ (and also close to $p$ up to such an additive error). So their Lemma 23 essentially states that estimating $p$ up-to additive $\ell_{2}$ error of $1/\sqrt{d}=\lambda\cdot\gamma$, requires $\tilde{\Omega}(\sqrt{d}/\lambda)$ users (which matches our Theorem 1.3 up to the low-order terms).
While both works yield similar smooth lower bound for DP averaging, the focus of the works is different: [NarayananME22] focus on user-level DP averaging, while our focus is on general bounds for various (item-level) DP problems. As part of that, we have several advantages in the item-level DP case:
@@ 2.4 Comparison with [NarayananME22]
    Tighter Bounds: [NarayananME22] have an $1/\log^{7}(d/\lambda)$ dependency in their DP-averaging lower bound (see their Theorem 10), while we only have an $1/\log^{1.5}(d/\lambda)$ dependency.
@@
    Boolean instances: We create boolean hard-instances, while [NarayananME22]’s instances are not.
@@ 4 Strong Fingerprinting Lemma
[…]
@@ 4.1 One Dimension
where we used the inequality $e^{-1.1y}\leq 1-y$ for every $y\in[0,1/6]$.
Hence, we conclude that
@@ 4.2 Increasing the Dimension
The first part of the lemma follows from similar arguments as in the proof of Proposition 10 in [DworkSSUV15]. In particular, for a given $j\in[d]$, we fix all columns except for the $j^{\rm th}$ column. Since the mechanism $\mathsf{M}$ only gets $\mathbb{X}$ as an input, and each columns $j$ of $\mathbb{X}$ is sampled independently, then there is some $F_{j}\colon\{-1,1\}^{n}\rightarrow[-1,1]^{d}$ for which $\mathbb{q}^{j}\sim F_{j}(\mathbb{x}^{j})$, where $\mathbb{x}^{j}=(\mathbb{x}_{1}^{j},\dots,\mathbb{x}_{n}^{j})$ (the $j$’th column of $\mathbb{X}$), such that Section 4.1 holds for it. Then, by this and the fact that $\mathsf{M}$ is a strongly-accurate mechanism,
@@
The second part is proven similarly to Lemma 11 in [DworkSSUV15]. Let us define the random variable $\mathbb{a}_{j}=\mathbb{q}^{j}\cdot\sum_{i\in[n]}(\mathbb{x}_{i}^{j}-\mathbb{z}^{j})$ for every $j\in[d]$. Then, we get that
@@
and from the first part we have ${\mathrm{E}}\mathopen{}\mathclose{{}\left[\sum_{j=1}^{d}\mathbb{a}_{j}}\right]\geq\frac{0.4d}{\ln(5n)}$. In order to use Hoeffding’s inequality on the sum $\sum_{j=1}^{d}\mathbb{a}_{j}$, we first assume that the random variables $\mathbb{a}_{1},\dots,\mathbb{a}_{d}$ are independent. Later, as in [DworkSSUV15], we show how to remove this assumption. Next, observe that $-2n\leq\mathopen{}\mathclose{{}\left|\mathbb{a}_{j}}\right|\leq 2n$ for every $j\in[n]$. Therefore, using an Hoeffding’s inequality (1), it follows that
@@ 5 Framework for Lower Bounds
Let $(\mathbb{x}_{1},\ldots,\mathbb{x}_{n_{0}},\mathbb{z})\sim{\cal D}(n_{0},d_{0})$, let $\mathbb{X}=(\mathbb{x}_{1},\ldots,\mathbb{x}_{n_{0}})$, and let $\mathbb{q}=\mathsf{A}(\mathbb{X})$.
By Section 5 and the union bound, we deduce that there exists $i\in[n_{0}]$ such that
@@
In the following, let $\mathbb{p}_{1},\ldots,\mathbb{p}_{d_{0}}\in[-1,1]$ be the (random variables of the) expectations that were chosen as part of the sampling of $(\mathbb{x}_{1},\ldots,\mathbb{x}_{n_{0}},\mathbb{z})$ from ${\cal D}(n_{0},d_{0})$ (as defined in Section 4.2), and in the rest of the proof we fix $(\mathbb{p}_{1},\ldots,\mathbb{p}_{d_{0}})=(p_{1},\ldots,p_{d_{0}})$ such that Equation 4 holds under this fixing.
For the above $i\in[n_{0}]$, let $\mathbb{x}_{i}^{\prime}\in\{-1,1\}^{d_{0}}$ be an independent random variable such that each $(\mathbb{x}_{i}^{\prime})^{j}\sim p_{j}$ (independently) for every $j\in[d_{0}]$. Note that $\mathbb{x}_{1},\ldots,\mathbb{x}_{n_{0}},\mathbb{x}_{i}^{\prime},\mathbb{z}$ are i.i.d. random variables. However, while $\mathbb{q}$ depends on $\mathbb{x}_{i}$, it is independent of $\mathbb{x}_{i}^{\prime}$ and $\mathbb{z}$.
Assume towards a contradiction that $\mathsf{A}$ is $\mathopen{}\mathclose{{}\left(1,\frac{\beta}{4n_{0}}}\right)$-DP. By applying Section 3.4.1 with the random predicate $P(x_{i},q)=\mathds{1}_{\{\langle x_{i},q\rangle-\langle\mathbb{z},q\rangle>\frac{0.2d_{0}}{n_{0}\ln(5n_{0})}\}}$, we deduce from Equation 4 that
@@
In the following, we prove that since $\mathbb{x}_{i}^{\prime}$ is independent of $\mathbb{q}$, then the above probability is much smaller, which will lead to a contradiction.
For every $j\in[d_{0}]$, define the random variable $\mathbb{b}_{j}=\mathbb{q}^{j}((\mathbb{x}_{i}^{\prime})^{j}-\mathbb{z}^{j})$. Then, we get that
@@ 6 Padding And Permuting (PAP) FPC
The first inequality hold by Section 6. The last inequality holds since $\mathsf{M}$ is $(\alpha,\beta)$-weakly-accurate and for every $X\in\{-1,1\}^{n\times d_{0}}$ and $Y\in\operatorname{Supp}(Y_{X})$ it holds that $\mathopen{}\mathclose{{}\left|{\cal J}_{Y}^{1}}\right|,\mathopen{}\mathclose{{}\left|{\cal J}_{Y}^{-1}}\right|\geq\ell\geq\frac{1}{2}(1-\alpha)d$.
Thus by Section 5, $n\geq\Omega(\sqrt{d_{0}}/\log^{1.5}(d_{0}/\beta))$, and the proof follows since $d_{0}=\Theta(\alpha d)$.     $\Box$
@@
We prove Theorem 1.2 using multiple PAP-FPC copies.
@@
    For $t\in[k]$ set $Y_{t}=\begin{cases}Y&t=s\\ \mathsf{PAP}(A_{t},P_{t})&t\neq s\end{cases}$.
    Denote $Y_{t}=(x_{(t-1)n_{0}+1}^{\prime},\>\ldots,\>x_{tn_{0}}^{\prime})\in\{-1,1\}^{n_{0}\times d}$.
@@
Define $\mathsf{G}\colon\{-1,1\}^{n_{0}\times d_{0}}\times{\cal V}\rightarrow\{-1,1\}^{n\times d}$ as the algorithm that on inputs $X\in\{-1,1\}^{n_{0}\times d_{0}}$ and $v=(s,\vec{P})\in{\cal V}$ for $s\in[k]$ and $\vec{P}=(P_{1},\ldots,P_{k})\in({\cal P}_{d})^{k}$: Computes $Y=\mathsf{PAP}(X,P_{s})$, and outputs $X^{\prime}\sim\mathsf{H}(Y,v)$. Note that for every $v\in{\cal V}$, $\mathsf{G}(\cdot,v)$ is neighboring-preserving (Section 3.4.1).
Define $\mathsf{F}\colon{\cal V}\times[-1,1]^{d}\rightarrow[-1,1]^{d_{0}}$ as the algorithm that on inputs $v=(s,\vec{P})\in{\cal V}$ and $w\in[-1,1]^{d}$, outputs $(w\cdot P_{s}^{T})^{1,\ldots,d_{0}}$. Note that $\mathsf{F}$ is deterministic, and the random choice of $\mathsf{G}$ is $\vec{A}\sim{\cal D}^{k}$ (chosen when executing $\mathsf{H}$).
Our goal is to show that
@@
We now focus on proving Equation 8. Consider the mechanism $\mathsf{M}^{\prime}\colon\{-1,1\}^{n_{0}\times d}\rightarrow[-1,1]^{d}$ that on input $Y\in\{-1,1\}^{n_{0}\times d}$, samples $v=(s,\vec{P})\leftarrow{\cal V}$, computes $X^{\prime}=\mathsf{H}(Y,v)$ and outputs $w\sim\mathsf{M}(X^{\prime})$.
In the following, let $\mathbb{P}\leftarrow{\cal P}_{k}$ be a random variable, and for $X\in\{-1,1\}^{n_{0}\times d_{0}}$ define $\mathbb{Y}_{X}=\mathsf{PAP}(X,\mathbb{P})$. Compute
@@
The first equality holds since by the definitions of $\mathsf{M}^{\prime}$ and $\mathsf{G}$, for every $X$, the pair $\mathopen{}\mathclose{{}\left(\mathsf{M}^{\prime}(\mathbb{Y}_{X}),\mathbb{Y}_{X}}\right)$ has the same distribution as the pair $\mathopen{}\mathclose{{}\left(\mathsf{M}(\mathsf{G}_{s,\vec{A}}(X,v)),\mathsf{PAP}(X,P_{s})}\right)_{v=(s,\vec{P})\leftarrow{\cal V},\>\vec{A}\sim{\cal D}^{k}}$. The second equality holds since $\mathsf{G}_{\vec{A}}(\cdot,(s,\vec{P}))$ does not use the value of $A_{s}$ which is drawn as $X$. The third equality holds by the definition of $\mathsf{G}$. The inequality holds since $\mathsf{M}$ is an $(k,\alpha,\beta)$-weakly-accurate mechanism, and it gets $X^{\prime}=(x_{1}^{\prime},\ldots,x_{n}^{\prime})$ where each $Y_{t}=(x^{\prime}_{(t-1)n_{0}+1},\>\ldots,\>x^{\prime}_{tn_{0}})$ has $\mathopen{}\mathclose{{}\left|{\cal J}_{Y_{t}}^{1}}\right|,\mathopen{}\mathclose{{}\left|{\cal J}_{Y_{t}}^{-1}}\right|\geq\ell\geq\frac{1}{2}(1-\alpha)d$. Thus w.p. $\beta$, $\mathsf{M}$ is guaranteed to output $w\in[-1,1]^{d}$ that strongly-agrees with one of the $Y_{1},\ldots,Y_{k}$. But since they are drawn independently from the same distribution, the probability to agree with $Y_{s}$ for a random $s\leftarrow[k]$ decreases by a factor $k$, and overall is at least $\beta/k$.
We now apply Section 6 to get that
@@ 7.1 Averaging
7.2 Clustering
In this section, we prove an extension of Theorem 1.4 to $(k,z)$-clustering where we focus on the Euclidean metric space $({\mathbb{R}}^{d},d(x,y):=\mathopen{}\mathclose{{}\left\lVert x-y}\right\rVert_{2})$. We start by extending the notations from Section 1.1.3.
Let ${\cal B}_{d}:=\{x\in{\mathbb{R}}^{d}\colon\mathopen{}\mathclose{{}\left\lVert x}\right\rVert_{2}\leq 1\}$. For a database $S\in({\cal B}_{d})^{n}$, $k$ centers $C=(c_{1},\ldots,c_{k})\in({\cal B}_{d})^{k}$ and a parameter $z\geq 1$, let
${\rm COST}_{z}(C;S):=\sum_{x\in S}\min_{i\in[k]}\mathopen{}\mathclose{{}\left\lVert x-c_{i}}\right\rVert_{2}^{z}\quad\text{ and }\quad{\rm OPT}_{k,z}(S):=\min_{C\in({\mathbb{R}}^{d})^{k}}{\rm COST}_{z}(C;S).$